Use the thermochemical data that follow to calculate the value of ΔH° for the reaction:
4 B 2 A + C + 2 F

A + 2B E ΔH° = 70 kJ
C + 2D 2E ΔH° = -25 kJ
F + 2A D ΔH° = 60 kJ
1
Expert's answer
2019-11-01T05:44:30-0400

Q97596


Solution:

As Enthalpy "\u0394H\u00b0" is an extensive property of a thermodynamic system, it changes with stoichiometric coefficients of the reaction.


Given:


"A + 2B \to E" "\u0394H\u00b0_1 = 70 kJ" ---(1)

"C + 2D \to 2E" "\u0394H\u00b0_2 = -25 kJ" ---(2)

"F + 2A \to D" "\u0394H\u00b0_3 = 60 kJ" ---(3)


To find: "4 B \to 2 A + C + 2 F" "\u0394H\u00b0=?"


Now, the required chemical equation can be obtained by doing the following;

"2*(1)-(2)-2*(3)"


Thus, the enthalpy of the required chemical reaction becomes "\u0394H\u00b0= 2\u0394H\u00b0_1- \u0394H\u00b0_2-2 \u0394H\u00b0_3"


"\u0394H\u00b0=(2*70)-(-25)-(2*60)"

"=140+25-120"

"=45kJ"